Solve the quadratic equations by completing the square.
x^2-6x-3=0
This one I do not have any idea.
Thank you for your help.

Answer by scott8148(6628)

completing the square means turning the quadratic into a "perfect square"

first, (using addition or subtraction) move the constant term to the other side of the equation

x^2-6x=3

next, (using multiplication or division) set the x^2 coefficient equal to 1 ... already done here

next, square 1/2 of the x term coefficient and add it to both sides ... -6/2=-3 ... (-3)^2=9

x^2-6x+9=12 ... (x-3)^2=12

next, take the square root of both sides ... remembering that square roots are + and -

x-3=sqrt(12) ... x-3=2(sqrt(3)) ... so x=3+2(sqrt(3)) and x=3-2(sqrt(3))